OpenStudy (dls):

Find the intervals in which f(x)=e^x^2 is increasing or decreasing

OpenStudy (dls):

I am getting (-inf,inf)

OpenStudy (dls):

\[\Huge f'(x)=e^{x^2} \times 2x \times 2 =4xe^{x^2}\] which is always greater than 0,since a^x>0 for all x

OpenStudy (shubhamsrg):

4x(e^x^2) is not always greater than 0

OpenStudy (shubhamsrg):

4(e^x^2) >0 always but not x my dear friend

OpenStudy (shubhamsrg):

inka to ho gaya ji good day! ^^

OpenStudy (dls):

haan,got it it can be equal to 0

OpenStudy (dls):

so root is only 0?

OpenStudy (shubhamsrg):

yo

OpenStudy (dls):

- + --------0----------

OpenStudy (dls):

cool

OpenStudy (dls):

medal do issi baat pe

OpenStudy (shubhamsrg):

maine bas 1 choti si chiz batayi us se itna zyada kaise samjh gaye? :O :O how! ? :O

OpenStudy (dls):

:")

OpenStudy (dls):

sarcasm tha?

OpenStudy (shubhamsrg):

nahi :| tum to mera nature jante hi ho, mazak to mai kabhi karta nahi

OpenStudy (dls):

bilkul mod ki tarah :)

OpenStudy (shubhamsrg):

90 SS pata nahi kaise mil gaya,! lol!!

OpenStudy (shubhamsrg):

and ye maine show off ke liye point out kiya :|

OpenStudy (dls):

ek aur question hai

OpenStudy (shubhamsrg):

pehli 1 ladki ki madad kar du :|